Some Ca(NO3)2 is dissolved in 100mL of water. This solution is mixed with 300mL of 0.010M Na2SO4. A very faint precipitate of calcium sulfate is formed. If the Ksp value of calcium sulfate is 2.4 x 10-5, how much calcium nitrate was dissolved to make the initial solution?
1
Expert's answer
2019-10-10T03:46:30-0400

The dissolution of calcium sulfate forms equal amounts of calcium ions and sulfate ions.

Then we can solve for the molarity of calcium sulfate that would be soluble.

2.4 x 105- = (x)*(x)

Solving for x we get [Ca2+] = [SO42-] = 4.90 x 10-3 M

Since the equation above shows a 1:1 mole ratio of calcium sulfate to Ca2+ ions, we can assume that 4.90 x 10-3 moles of CaSO4 will dissolve.

Na2SO4 + Ca(NO3)2 = CaSO4 + 2NaNO3

n(Na2SO4) = C*V = 0.01mol/L*0.3L =0.003 mol

According to the equation 0.003 mol of Na2SO4 reacts with 0.003 mol of Ca(NO3)2, and this amount corresponds to that specified in Ksp .

n(Ca(NO3)2) = 0.003 mol

m(Ca(NO3)2) = n*M = 0.003mol*164g/mol = 0.492 g